OpenStudy (anonymous):

The function H(t) = -16t2 + vt + s shows the height H (t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 100 feet with an initial velocity of 60 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. (2 points) Part B: What is the maximum height that the projectile will reach? Show your work. (2 points) Part C: Another object moves in the air along the path of g(t) = 20 + 38.7t where g(t

OpenStudy (anonymous):

he height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know? (2 points)

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

I think for part A, the acceleration is the second derivative of position. Maximum height occurs when the instantanous velocity is zero. So you take the first derivative and equate it to zero. The third part is incomplete.

OpenStudy (anonymous):

Part C: Another object moves in the air along the path of g(t) = 20 + 38.7t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points)

OpenStudy (anonymous):

Could you explain part A?

OpenStudy (anonymous):

Is this question in physics or math? Calculus one?

OpenStudy (anonymous):

math, algebra 1

OpenStudy (anonymous):

Oh..... I way over thought this problem sorry.

OpenStudy (anonymous):

hahahaha its okay. i don't understand it at all so.

OpenStudy (anonymous):

For the first part, it is stated that the height is 100 feet and initial velocity is 60 feet per second. So, in order to get to time, you can plug these values in the formula. h(t) = 100 and v = 60. Can you see what I am talking about?

OpenStudy (anonymous):

whats the formula?

OpenStudy (anonymous):

Its given. The first one on top.

OpenStudy (anonymous):

oh wait never mind, let me try it.

OpenStudy (anonymous):

So it will be: 100 = -16t^2 +60t + s. So far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you solve for t?

OpenStudy (anonymous):

wouldn't you need s?

OpenStudy (anonymous):

Its not given, dont worry about it.

OpenStudy (anonymous):

okay so it would be

OpenStudy (anonymous):

Okay good. Now, for part B, the total time to get there is half of this so you take this value and divide it by 2. After doing that you take and put it the equation you got for part A for t, and then you solve for s. This will give you the maximum height reached.

OpenStudy (anonymous):

No the value of the t divided by 2. You plug that value into this formula and you solve for s.